OpenStudy (anonymous):

I'll give a medal again please help!! Juliana has created the function f(x) = the quantity of 3x plus 2, divided by 4 to represent the cost of texting on her current plan, where x represents the number of texts. Juliana discovers that, using the inverse function to solve for x = 24, she can predict how many texts she can use for $24. Explain to Juliana how to accomplish this, using complete sentences.

OpenStudy (anonymous):

The equation that Juliana created is \[f(x)=\frac{ 3x+2 }{ 4 }\]. Plug in 24 for x. She wants to be able to know the amount of texts she can send with $24. Solve the equation by dividing.

OpenStudy (anonymous):

so it would 3x+2= 4 right?

OpenStudy (anonymous):

So I just realized that your problem says find the inverse function.. Okay, to find the inverse of the function, you need to set f(x) into y and exchange the positions of y and x in the equation then solve for y. \[f(x)=\frac{ 3x+2 }{ 4 }\] would turn into \[y=\frac{ 3x+2 }{ 4 }\] \[x=\frac{ 3y+2 }{ 4 }\] From here, isolate the y value.

OpenStudy (anonymous):

By isolating the value, you need to put 3y on one side of the equation.

OpenStudy (anonymous):

so 3y+2 over 4

OpenStudy (anonymous):

Work from \[x=\frac{ 3y+2 }{ 4 }\] Isolate the y by multiplying both sides of the equation with 4. By multiplying both sides of the equation by 4, it will cancel out dividing the 4 on the y side.

OpenStudy (anonymous):

After you've multiplied both sides by 4, the equation is now 4x = 3y + 2. You have to subtract 2 from both sides of the equation again to isolate y.

OpenStudy (anonymous):

do i subtract the 2 from the y or x

OpenStudy (anonymous):

You subtract them from both.

OpenStudy (anonymous):

You have to get y by itself. After you've subtract 2 from both sides, 3y is still on one side of the equation. How would you get 3y to turn into y?

OpenStudy (anonymous):

by multiply the 3y right

OpenStudy (anonymous):

No, you divide the 3y to get 3. Now, the equation is \[y = \frac{ 4x+2 }{ 3y }\] That's the inverse function of the original function that Juliana created. Now, plug in 24 for x.

OpenStudy (anonymous):

"\[x=\frac{4x-2 }{ 3 }\]" sorry my mistake

OpenStudy (anonymous):

Now what do i do?

OpenStudy (anonymous):

Do you want to this in private chat?

OpenStudy (anonymous):

Plug in 24 for x. That's the last calculation

OpenStudy (anonymous):

so 24x-2 over 3

OpenStudy (anonymous):

\[x=\frac{4(24)-2 }{ 3 }\]

OpenStudy (anonymous):

ok i check that and divide it i got a decimal , is that fine?

OpenStudy (anonymous):

You possibly can't have a decimal as the answer. Julianna can't send a half of a text. You need to calculate the numerator first. Then, divide the numerator by the denominator.

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

But there's still a decimal

OpenStudy (anonymous):

Well, try to do the first part of the actual equation. What is 4 * 24?

OpenStudy (anonymous):

ok, that is 96

OpenStudy (anonymous):

Subtract 2 from 96.

OpenStudy (anonymous):

-94

OpenStudy (anonymous):

96 - 2 is 94. Divide by 3. Round your answer so that it doesn't have any decimals.

OpenStudy (anonymous):

31

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

did you get your medal?